# Overcoming Long-Context Limitations of State-Space Models via Context-Dependent Sparse Attention

### v2

@@ 5.1 Empirical Verification on Multi-Query Joint Recall
Baselines. We adopt Mamba (Gu & Dao, 2024) and Mamba2 (Dao & Gu, 2024) as base architectures and evaluate various hybrid sparse attention models built upon them, as illustrated in Figure 3. These include dilated attention (D), sliding window attention (SW), a combination of sliding window and dilated attention (SW+D) (Ding et al., 2023a), A-shaped attention (A), locality-sensitive hashing (LSH) attention, our proposed key selection (KS) attention, and HAX (a combination of LSH and KS attention). We also consider a Samba baseline. For fair comparison, we double the hidden size of SSMs, and fix $k$ defined in Eq. 4 (the maximum number of keys each query can attend to) across all hybrid architectures.
Setup. For evaluation, we calculate mean accuracy per-sample and average over the testset. For each base architecture and hybrid sparse attention variant, we conduct experiments using $3$ different random seeds for each learning rate in $\{3$e-33, 11e-33, 33e-4}4\}, and report the average performance corresponding to the best-performing learning rate.
Results. The results are summarized in Table 5. Compared to the Mamba and Mamba2 base architectures, most hybrid sparse attention models show performance improvements. In particular, our proposed HAX model consistently achieves the best performance, which underscores the effectiveness of our approach.
@@ 5.2 Continual Pre-training on Natural Language
Setup. To evaluate our method on real-world long-context natural language data, we perform continual pre-training based on the publicly released Mamba checkpoints (Gu & Dao, 2024). As in section 5.1, we augment the Mamba architecture with sparse attention, as illustrated in Figure 3. We include SSMs integrated with CISA as baselines, along with ablated variants of our HAX model. Overall, our HAX adds only about 1% additional parameters to the released Mamba architecture. All models are continually pre-trained for 50K steps with a context length of 2K.
Validation Loss during Continual Pre-Training. Figure 5 shows the validation loss $\mathcal{L}_{LM}$ of 130M model continual pre-training on The Pile (Gao et al., 2021). As observed, the Mamba base architecture and all baseline variants exhibit either training instability or plateau early in the training process. In contrast, our proposed HAX model is the only architecture that shows a consistent decline in validation loss throughout the training process, indicating improved stability and sustained learning.
Ruler and LongBench Evaluation. Tables 1 and 2 present the downstream performance for the Mamba 790M model after continual pre-training for 50K steps on TxT360 (Tang et al., 2024) followed by instruction fine-tuning for 10K steps on UltraChat (Ding et al., 2023b) following Mamba-Chat (Mattern & Hohr, 2023). Ruler (Hsieh et al., 2024) is a synthetic long-context NLP benchmark designed to assess model performance on tasks including retrieval, multi-hop reasoning, aggregation, and question answering. LongBench (Bai et al., 2024) comprises real-world NLP long-context tasks, including question answering, summarization, few-shot learning, retrieval, aggregation, and code completion. The evaluation results, summarized in Tables 1 and 2, show that among all hybrid sparse attention variants, our proposed HAX model is the only one that outperforms the Mamba baseline by a significant margin on average.
@@ 6.1 State-Space Models
State-space models (SSMs) originated in control theory, exemplified by damped mass-spring systems (Patro & Agneeswaran, 2024). HiPPO (Gu et al., 2020) was one of the first efforts to adapt SSMs for machine learning applications. LSSL (Gu et al., 2021) unified convolutional neural networks (CNNs), recurrent neural networks (RNNs), and ordinary differential equations (ODEs) under the SSM framework, enabling their implementation within deep neural networks. H3 (Fu et al., 2023) integrated SSM layers with short convolutional filters to enhance sequence modeling. Mamba (Gu & Dao, 2024) advanced this line of work by making all SSM parameters input-dependent, which significantly increases the representation capacity of SSMs. Mamba2 (Dao & Gu, 2024) further improved the architecture and established connections between SSMs and Transformer attention.
@@ 6.2 Analysis on State-Space Models
Recent empirical studies have shown that SSM long-context modeling performance often lags behind that of Transformer architectures (Waleffe et al., 2024). Jelassi et al. (2024) demonstrated that even solving simple tasks like copying requires SSM state dimensions to grow linearly with the sequence length. Furthermore, Merrill et al. (2024) established that the expressiveness of SSMs is bounded by the complexity class TC0. Sarrof et al. (2024) further showed that SSMs and Transformers capture overlapping yet distinct subsets of TC0, providing a theoretical basis for developing hybrid models that combine the strengths of both architectures.
@@ 6.3 Hybrid Architectures
Several works have explored architectures that mix a large proportion of SSM layers with a small number of full attention layers, and have reported performance surpassing that of standard Transformers (Waleffe et al., 2024). The effectiveness of such hybrid architectures has been further validated at billion-parameter scale (Lenz et al., 2025). In parallel, researchers have also investigated the design of hybrid sparse attention models (Ren et al., 2025; Dong et al., 2025; Nunez et al., 2024), which offer sub-quadratic computational complexity, providing a promising direction for efficient long-context modeling.
@@ Appendix B Theoretical Proof
As a direct consequence of Theorem 2.7 in Jelassi et al. (2024), we adopt their proof strategy. We reformulate Lemma D.1 in Jelassi et al. (2024) as the following Lemma B.1. Let $m$ denote the index of the last token in the information component. Then, for any fixed permutation $\mathcal{P}$ of the context and keys in the inquiry component, the following Lemma B.1 holds:
@@
Following Jelassi et al. (2024), we bound the error of the model by comparing the number of possible model states to the number of distinct input instances.
@@
In contrast, as illustrated in Fig. 7, hybrid sparse attention models mitigate this limitation by enabling information to propagate through multiple parallel paths, thereby alleviating the bottleneck imposed by sequential state updates.
@@ C.2 Expressiveness of SSMs Integrated with CDSA on Multi-Level Context Joint Recall
Table 3: Results of Ruler benchmark at 4K context length. We compare different sparse attention integrated with Mamba, including CISA and CDSA methods as in Tab. 1. The best average performance is in bold.
Mamba 100 19.2 9.2 5.2 0.6 0 0.65 1.85 5.72 4.3 100 26.2 78.6 27.04
+D 0 0 0 0 0 0 0 0 65.2 0 100 0 73 18.32
+SW 100 18.2 5 5.6 0.2 0.2 1.75 2.1 4.4 4.78 100 22.4 78.8 26.42
+SW+D 100 20.2 8.2 6.2 0.4 0.2 1.45 1.25 4.36 4.32 100 24.6 79 26.94
+A 100 19.8 7.2 6.6 1.2 0.2 1.3 0.8 4.36 3.24 100 24.2 79 26.76
+LSH 0 0 0 0 0 0 0 0 0 0.04 100 0.2 73 13.33
+KS (ours) 99.8 20.2 7.8 8.2 1 0.2 1.2 1.85 9.32 2.86 100 22.6 78.6 27.20
+HAX (ours) 100 23 11 9 0.8 0 3.4 2.75 6.36 2.54 100 29 78.8 28.20
@@ D.1 Empirical Verification on Joint Recall
For all models, we fix the number of layers to 2, set the hidden size to 64, and use $k=64$. For variants that integrate multiple sparse attention components, namely SW+D, A (consists of a SW component and a sink attention component, a sink attention always attends to the first $k$ tokens in the sequence only), and HAX (LSH+KS), we allocate $k=32$ to each component, in order to maintain a global $k=64$. For both LSH and HAX (LSH+KS), we adopt the sign-bit binning strategy (Eq. 8) with $h=8$, and refresh the random hashing matrix at each training step. To account for the absence of sparse attention mechanisms, we additionally evaluate Mamba and Mamba2 baselines with their hidden size doubled to 128, ensuring a fair comparison. Additionally, we include a Samba baseline, consisting of 2 Mamba layers and 2 sliding window attention layers. For Samba, the hidden size and sliding window width $(k)$ are both set to 64. We use AdamW optimizer. All models are trained for 400,000 steps with a batch size of 64. Our implementation is based on Flash-Linear-Attention (Yang & Zhang, 2024).
@@ D.2 Continual Pre-training on Natural Language
For all experiments, we fix the sparsity parameter at $k=128$. For variants that integrate multiple sparse attention components, namely SW+D, A, and HAX (LSH+KS), we set $k=64$ to each component, in order to maintain a global $k=128$. For both LSH and HAX (LSH+KS), we adopt the argmax binning strategy (Eq. 7) with $h=k$. We resample the random hashing matrix at each training step and fix a random hashing matrix during evaluation.
To enhance the long-context modeling capability, we filter samples to retain only those with tokenized lengths of at least 2,048 tokens. At the beginning of continual pre-training, the $\mathbf{K}$} and $\mathbf{Q}$} projection weights are initialized using the parameters of the $\mathbf{B}$} and $\mathbf{C}$} projections from Mamba, respectively, based on state-space duality (Dao & Gu, 2024). During continual pretraining, we use a cosine schedule with a maximum learning rate of 3e-4 for 130M models and 1.5e-4 for 790M models, and a minimum learning rate of 1e-5. A warm-up phase of 200 steps with a learning rate of 0 followed by 800 steps of linear warm-up precedes the cosine schedule. For instruction tuning, we also apply a 200-step 0 learning rate phase followed by 800 steps of linear warm-up, after which the learning rate remains constant at 3e-6. For both continual pre-training and instruction-tuning, we use AdamW optimizer and train with a global batch size of 64 and a context length of 2K.
@@ Appendix E Additional Experiments
E.1 Extrapolation Results on the Ruler Benchmark at 4K Context Length
We additionally evaluate all models on the Ruler (Hsieh et al., 2024) benchmark at a 4K context length, despite being trained solely with a 2K context length. As summarized in Table 3, our proposed HAX model consistently outperforms all baselines, underscoring its robustness for context length extrapolation.
E.2 Short-context Modeling Benchmark
We follow Mamba (Gu & Dao, 2024) to evaluate the zero-shot short context modeling performance of the continually pre-trained models on the LM evaluation harness benchmark from EleutherAI (Gao et al., 2024). Our results in Table4 show that continual pre-training on long sequences will not lead to a significant performance drop on short context benchmarks, where the Mamba w/o continual pre-training results are copied from the Mamba paper (Gu & Dao, 2024).
Table 4: LM evaluation harness benchmark for continually pre-trained Mamba models. We compare different sparse attention integrated with Mamba, including CISA and CDSA methods as in Tab. 1.
@@ E.2 Short-context Modeling Benchmark
Figure 10: An example of the "vertical-stripe" attention pattern in LLM. We input the first paragraph of the Wikipedia term "Harry Potter" (Wikipedia, 2025) into the Llama 3.2 1B model (Grattafiori et al., 2024) and visualize the log attention probabilities of the last head in the first layer. The input text is: “Harry Potter is a series of seven fantasy novels written by British author J. K. Rowling. The novels chronicle the lives of a young wizard, Harry Potter, and his friends, Ron Weasley and Hermione Granger, all of whom are students at Hogwarts School of Witchcraft and Wizardry. The main story arc concerns Harry’s conflict with Lord Voldemort, a dark wizard who intends to become immortal, overthrow the wizard governing body known as the Ministry of Magic, and subjugate all wizards and Muggles (non-magical people).”

### v3

@@ 5.1 Empirical Verification on Multi-Query Joint Recall
Baselines. We adopt Mamba mamba and Mamba2 mamba2 as base architectures and evaluate various hybrid sparse attention models built upon them, as illustrated in Figure 3. These include dilated attention (D), sliding window attention (SW), a combination of sliding window and dilated attention (SW+D) longnet , A-shaped attention (A), locality-sensitive hashing (LSH) attention, our proposed key selection (KS) attention, and HAX (a combination of LSH and KS attention). We also consider a Samba baseline.
Setup. For fair comparison, we fix the hidden size to $64$ and fix $k=64$ defined in Eq. 4 (the maximum number of keys each query can attend to) across all hybrid architectures. For variants that integrate multiple sparse attention components, namely SW+D, A (consists of a SW component and a sink attention component, a sink attention only attends to the first tokens in the sequence), and HAX (LSH+KS), we allocate $k=32$ to each component, in order to maintain a global $k=64$. We also consider a baseline that doubles the hidden size of SSMs to compensate for the absence of hybrid sparse attention. We fix the number of layers to $2$ for all models. We conduct experiments using $3$ different random seeds for each learning rate in $\{3$e-33, 11e-33, 33e-4}4\}, and report the average performance corresponding to the best-performing learning rate. For evaluation, we calculate mean accuracy per-sample and average over the testset.
Results. As summarized in Table 5, compared to the Mamba and Mamba2 base architectures, most hybrid sparse attention models improve performance. In particular, our proposed HAX model consistently achieves the best performance, which underscores the effectiveness of our approach.
Table 1: Ruler benchmark at 2K context length. We compare different sparse attention integrated with Mamba 790M model, including CISA methods: dilated attention (D), sliding window attention (SW), and their combination (SW+D), and A-shaped attention (A), and CDSA methods: LSH attention, and our proposed key selection (KS) attention and HAX. The best average performance is in bold.
Mamba 100 95.2 86.8 34.4 2.6 0 32.4 28.35 2.76 1.4 100 25.2 42.43
+D 0 0 0 0 0 0 0 0 17.4 0 100 0 9.78
+SW 100 93.6 83.6 36.2 3.4 0 31.05 31.45 5.24 1.26 100 24.4 42.52
+SW+D 100 95 82.4 34 2.6 0 32.9 29.45 4.72 1.22 100 24.8 42.26
+A 100 93.2 81.4 34.4 2.6 0 31.7 27.1 4.4 1.26 100 26.8 41.91
+LSH 0 0 0 0 0 0 0 0 0.2 0.06 100 0 8.36
+KS (ours) 100 94.4 84.4 33.6 1.8 1.8 37.25 33.85 4.04 1 100 25 43.10
+HAX (ours) 100 98.8 89.6 45.4 3.8 0.2 35.5 28.2 2.04 1.64 100 29.6 44.57
Table 2: LongBench English tasks. We compare different sparse attention integrated with Mamba 790M model, including CISA and CDSA methods. The best average performance is in bold.
+D 0 0.21 0 2.87 0.26 0 0 0 0 0 0 0.72 3.46 0.7 0 0 0.51
Table 3: Performance of extrapolation from 2K to 4K context length on Ruler benchmark. We compare different sparse attention integrated with Mamba 790M model, including CISA and CDSA methods as in Table 3. The best average performance is in bold.
Mamba 100 19.2 9.2 5.2 0.6 0 0.65 1.85 5.72 4.3 100 26.2 78.6 27.04
+D 0 0 0 0 0 0 0 0 65.2 0 100 0 73 18.32
+SW 100 18.2 5 5.6 0.2 0.2 1.75 2.1 4.4 4.78 100 22.4 78.8 26.42
+SW+D 100 20.2 8.2 6.2 0.4 0.2 1.45 1.25 4.36 4.32 100 24.6 79 26.94
+A 100 19.8 7.2 6.6 1.2 0.2 1.3 0.8 4.36 3.24 100 24.2 79 26.76
+LSH 0 0 0 0 0 0 0 0 0 0.04 100 0.2 73 13.33
+KS (ours) 99.8 20.2 7.8 8.2 1 0.2 1.2 1.85 9.32 2.86 100 22.6 78.6 27.20
+HAX (ours) 100 23 11 9 0.8 0 3.4 2.75 6.36 2.54 100 29 78.8 28.20
@@ 5.2 Continual Pre-training on Natural Language
Setup. To evaluate our method on real-world long-context natural language data, we perform continual pre-training based on the publicly released Mamba checkpoints mamba . As in section 5.1, we augment the Mamba architecture with sparse attention, as illustrated in Figure 3. We include SSMs integrated with CISA as baselines, along with ablated variants of our HAX model. We fix the sparsity parameter $k=128$. For variants that integrate multiple sparse attention components, namely SW+D, A, and HAX (LSH+KS), we set $k=64$ to each component, in order to maintain a global $k=128$. Overall, our HAX adds only about 1% additional parameters to the released Mamba architecture. All models are continually pre-trained for 50K steps with a context length of 2K.
Validation Loss during Continual Pre-Training. Figure 5 shows the validation loss $\mathcal{L}_{LM}$ of 130M model continual pre-training on The Pile pile . As observed, the Mamba base architecture and all baseline variants exhibit either training instability or plateau early in the training process. In contrast, our proposed HAX model is the only architecture that shows a consistent decline in validation loss throughout the training process, indicating improved stability and sustained learning.
Ruler and LongBench Evaluation. Tables 3 and 3 present the downstream performance for the Mamba 790M model after continual pre-training for 50K steps on TxT-360 txt360 followed by instruction fine-tuning for 10K steps on UltraChat ding2023enhancing following Mamba-Chat haven2023mambachat . Ruler ruler is a synthetic long-context NLP benchmark designed to assess model performance on tasks including retrieval, multi-hop reasoning, aggregation, and question answering. LongBench longbench comprises real-world NLP long-context tasks, including question answering, summarization, few-shot learning, retrieval, aggregation, and code completion. The evaluation results, summarized in Tables 3 and 3, show that among all hybrid sparse attention variants, our proposed HAX model is the only one that outperforms the Mamba baseline by a significant margin on average.
Extrapolation Evaluation. We additionally evaluate the Mamba 790M models on the Ruler ruler benchmark at a 4K context length, despite being continually pre-trained and instruction fine-tuned solely with a 2K context length. As summarized in Table 3, our proposed HAX model consistently outperforms all baselines, underscoring its robustness for context length extrapolation.
@@ 6.1 State-Space Models
State-space models (SSMs) originated in control theory, exemplified by damped mass-spring systems patro2024mamba360 . HiPPO hippo was one of the first efforts to adapt SSMs for machine learning applications. LSSL lssl unified convolutional neural networks (CNNs), recurrent neural networks (RNNs), and ordinary differential equations (ODEs) under the SSM framework, enabling their implementation within deep neural networks. H3 h3 integrated SSM layers with short convolutional filters to enhance sequence modeling. Mamba mamba advanced this line of work by making all SSM parameters input-dependent, which significantly increases the representation capacity of SSMs. Mamba2 mamba2 further improved the architecture and established connections between SSMs and Transformer attention.
@@ 6.2 Analysis on State-Space Models
Recent empirical studies have shown that SSM long-context modeling performance often lags behind that of Transformer architectures waleffe2024empirical . jelassi2024repeat demonstrated that even solving simple tasks like copying requires SSM state dimensions to grow linearly with the sequence length. Furthermore, ssmillusion established that the expressiveness of SSMs is bounded by the complexity class TC0. sarrof2024formallanguage further showed that SSMs and Transformers capture overlapping yet distinct subsets of TC0, providing a theoretical basis for developing hybrid models that combine the strengths of both architectures.
6.3 Context-Dependent Sparse Attention
Locality-sensitive hashing (LSH) attention reformer stands as one of the most widely adopted forms of context-dependent sparse attention. Follow-up LSH attention variants can be broadly categorized by their binning rules: the argmax binning rule is adopted in reformer ; lha , while the sign-bit binning rule is used in memoryformer ; hashatt . scfa introduces a Triton kernel that accelerates hash-based sparse attention. Recently, nsa proposed native sparse attention, a new context-dependent sparse attention that outperforms Transformers training from scratch, highlighting its potential for efficient long-context modeling.
@@ 6.4 Hybrid Architectures
Several works have explored architectures that mix a large proportion of SSM layers with a small number of full attention layers, and have reported performance surpassing that of standard Transformers waleffe2024empirical . The effectiveness of such hybrid architectures has been further validated at billion-parameter scale jamba . In parallel, researchers have also investigated the design of hybrid sparse attention models sma ; samba ; hymba ; nunez2024expansionspan ; bmojo , which offer sub-quadratic computational complexity, providing a promising direction for efficient long-context modeling.
@@ 7 Conclusion
Acknowledge
We would like to thank Jingyang Yuan, Wanru Zhao, Meng Qu, Zewen Chi for their insightful discussion on language model pre-training. We also thank the anonymous reviewers, AC, SAC, and PC for their time, thoughtful feedback, and constructive engagement throughout the review process.
@@ Appendix B Theoretical Proof
As a direct consequence of Theorem 2.7 in [19], we adopt their proof strategy. We reformulate Lemma D.1 in [19] as the following Lemma B.1. Let $m$ denote the index of the last token in the information component. Then, for any fixed permutation $\mathcal{P}$ of the context and keys in the inquiry component, the following Lemma B.1 holds:
@@
Following [19], we bound the error of the model by comparing the number of possible model states to the number of distinct input instances.
@@
In contrast, as illustrated in Figure 7, hybrid sparse attention models mitigate this limitation by enabling information to propagate through multiple parallel paths, thereby alleviating the bottleneck imposed by sequential state updates.
@@ D.1 Empirical Verification on Joint Recall
For all models, we fix the number of layers to 2, set the hidden size to 64, and use $k=64$. For variants that integrate multiple sparse attention components, namely SW+D, A (consists of a SW component and a sink attention component, a sink attention always attends to the first $k$ tokens in the sequence only), and HAX (LSH+KS), we allocate $k=32$ to each component, in order to maintain a global $k=64$. For both LSH and HAX (LSH+KS), we adopt the sign-bit binning strategy (Eq. 7) with $h=8$, and refresh the random hashing matrix at each training step. To account for the absence of sparse attention mechanisms, we additionally evaluate Mamba and Mamba2 baselines with their hidden size doubled to 128, ensuring a fair comparison. Additionally, we include a Samba baseline, consisting of 2 Mamba layers and 2 sliding window attention layers. For Samba, the hidden size and sliding window width $(k)$ are both set to 64. We use AdamW optimizer. All models are trained for 400,000 steps with a batch size of 64. Our implementation is based on Flash-Linear-Attention [42].
@@ D.2 Continual Pre-training on Natural Language
For all experiments, we fix the sparsity parameter $k=128$. For variants that integrate multiple sparse attention components, namely SW+D, A, and HAX (LSH+KS), we set $k=64$ to each component, in order to maintain a global $k=128$. For both LSH and HAX (LSH+KS), we adopt the argmax binning strategy (Eq. 6) with $h=k$. We resample the random hashing matrix at each training step and fix a random hashing matrix during evaluation.
To enhance the long-context modeling capability, we filter samples to retain only those with tokenized lengths of at least 2,048 tokens. At the beginning of continual pre-training, the $\mathbf{K}$} and $\mathbf{Q}$} projection weights are initialized using the parameters of the $\mathbf{B}$} and $\mathbf{C}$} projections from Mamba, respectively, based on state-space duality [4]. During continual pretraining, we use a cosine schedule with a maximum learning rate of 3e-4 for 130M models and 1.5e-4 for 790M models, and a minimum learning rate of 1e-5. A warm-up phase of 200 steps with a learning rate of 0 followed by 800 steps of linear warm-up precedes the cosine schedule. For instruction tuning, we also apply a 200-step 0 learning rate phase followed by 800 steps of linear warm-up, after which the learning rate remains constant at 3e-6. For both continual pre-training and instruction-tuning, we use AdamW optimizer and train with a global batch size of 64 and a context length of 2K.
@@ Appendix E Additional Experiments
E.1 Short-context Modeling Benchmark
We follow Mamba [14] to evaluate the zero-shot short context modeling performance of the continually pre-trained models on the LM evaluation harness benchmark from EleutherAI [12]. Our results in Table 4 show that continual pre-training on long sequences will not lead to a significant performance drop on short context benchmarks, where the Mamba w/o continual pre-training results are copied from the Mamba paper [14].
Table 4: LM evaluation harness benchmark for continually pre-trained Mamba models. We compare different sparse attention integrated with Mamba, including CISA and CDSA methods as in Table 3.
@@ E.1 Short-context Modeling Benchmark
Figure 10: An example of the "vertical-stripe" attention pattern in LLM. We input the first paragraph of the Wikipedia term "Harry Potter" [41] into the Llama 3.2 1B model [13] and visualize the log attention probabilities of the last head in the first layer. The input text is: “Harry Potter is a series of seven fantasy novels written by British author J. K. Rowling. The novels chronicle the lives of a young wizard, Harry Potter, and his friends, Ron Weasley and Hermione Granger, all of whom are students at Hogwarts School of Witchcraft and Wizardry. The main story arc concerns Harry’s conflict with Lord Voldemort, a dark wizard who intends to become immortal, overthrow the wizard governing body known as the Ministry of Magic, and subjugate all wizards and Muggles (non-magical people).”